# CryoSAMU: Enhancing 3D Cryo-EM Density Maps of Protein Structures at Intermediate Resolution with Structure-Aware Multimodal U-Nets

### v1

@@ 3.1 Dataset of protein structures and density maps
Our dataset was built with a set of cryo-EM density maps at resolutions from 4.0 Å to 7.9 Å from the EMDB databank [24] and their associated protein structures from the PDB databank [25]. To ensure that density maps are properly aligned with their corresponding PDB structures, we excluded maps and PDBs from the dataset if: (i) maps contain extensive regions without or misaligned corresponding PDB structures; (ii) maps contain other macromolecules except proteins; (iii) PDB structures contain backbone atoms only and/or unknown residues. Furthermore, to enhance training efficiency, we measured the correlation between map-PDB pairs using ChimeraX [26], and removed pairs with correlation score lower than 0.65. To avoid data redundancy, we measured the sequence identity between PDB structures, and retained only one if identity is greater than 30 %. As a result, a total of 384 pairs of cryo-EM maps and associated PDB structures remained. Among these data, 247 ($\sim$65 %), 62 ($\sim$15 %), and 75 ($\sim$20 %) map-PDB pairs were selected as training, validation, and test sets, respectively. Details are listed in Supplementary Tables 1-3.
@@ 4.1 Visualization and quantification of map enhancement
As displayed in Figure 2b, CryoSAMU significantly suppressed noise in the lip nanodisc regions (highlighted by dashed boxes in Figure 2a) of the deposited map for EMDB-24496 (PDB-7RKF). Moreover, the deposited map at a smaller volumes missed certain structural regions corresponding to the protein structures, as highlighted by black boxes in Figure 2a. In contrast, the CryoSAMU-enhanced maps exhibited better alignment with the corresponding protein structures, revealing more structural details, as demonstrated by black boxes in Figure 2b. Similar visual results were observed for another protein structure (see Supplementary Figure 1). Furthermore, residue-level real-space correlation coefficient (RSCC) measurements [35] in Figure 2c suggested significant improvements. Specifically, Chains A, B, and C exhibit RSCC increases compared to the deposited map, with correlations rising from 0.835 to 0.860, 0.848 to 0.865, and 0.801 to 0.836, respectively. In addition, 84.9%, 73.5%, and 90.6% of resiudes in Chains A, B, and C, respectively, showcased higher RSCC scores. Consistent RSCC improvements were also observed in other samples (see Supplementary Figure 2).
@@ 4.2 Benchmark I: improvement of real and Fourier space correlations
CC_box 0.731 0.679 0.618 0.862 0.834 0.751
CC_peaks 0.750 0.722 0.611 0.774 0.753 0.698
CC_volume 0.594 0.542 0.534 0.729 0.691 0.571
FSC05 6.124 6.147 5.283 4.668 5.108 6.434
@@ 4.3 Benchmark II: improvement of protein structure modeling
As the goal of enhancing cryo-EM density maps is to improve the performance of protein structure modeling from density maps (i.e., map interpretability), we benchmarked protein structures constructed from CryoSAMU-enhanced maps against those processed by other methods. Specifically, we used a standard structure modeling tool, known as phenix.map_to_model [36], to construct protein structures from 20 maps enhanced by the different tested methods. These maps were randomly selected from the test dataset to ensure that they were not exposed during training, as listed in Supplementary Table 4. To evaluate these structures, we used phenix.chain_comparison [36] to compare the constructed structures against their corresponding ground-truth PDB protein structures. We reported two metrics: residue coverage and sequence match. The residue coverage indicates the fraction of residues in the query structure that match the corresponding residues in the target structure within 3.0 Å, regardless of residue type. The sequence match indicates the percentage of matched residues that share identical residue types between the query and target structures.
@@ 5 Conclusion and Discussion
In this work, we introduce CryoSAMU, the first structure-aware multimodal network for enhancing cryo-EM density maps at intermediate resolution of protein structures. Our approach combines 3D map features with corresponding structural features through cross-attention mechanisms. In addition, we develop a self-attention weighting algorithm to produce fixed-size representations of structural embeddings derived from the pretrained ESM-IF1 model, preserving inter-chain and residue relationships while maintaining structural integrity. Our benchmark results demonstrate that CryoSAMU preforms competitively with existing cutting-edge methods, closely approaching the performance of EMReady, the current leading tool for cryo-EM density map enhancement. Notably, CryoSAMU achieves the fastest processing speed among all tested methods, positioning it as a promising solution for large-scale and practical applications. Furthermore, our ablation study reveals that incorporating an additional structural modality significantly boosts CryoSAMU’s performance across all evaluation metrics, suggesting a new avenue for future cryo-EM research to explore the effective integration of multimodal data during network training.
@@ Acknowledgments
Figure 1: Visualizations of deposited (blue) and CryoSAMU-enhanced (green) maps. The corresponding PDB structures (brown) are superimposed on the maps. a Human Dispatched-1 (PDB-6XE6, EMDB-22144, reported resolution of 4.53 Å) [37]. a-b: Maps displayed at the recommended contour level. c-d: Maps displayed at a higher contour level. Visualizations were produced by UCSF ChimeraX [26]. The protein structure modeling completeness and accuracy improved after CryoSAMU enhancement. For instance, residue coverage increased from 53.5% to 64.5%, as well as sequence match increased from 6.1% to 7.7%.
Figure 2: The real-space correlation coefficient (RSCC) comparison between deposited and CryoSAMU-enhanced maps. Top: a transmembrane nanopore TMH4C4 (PDB-6M6Z, EMDB-30126, reported resolution of 5.9 Å) [38]. Bottom: a yeast cytoplasmic exosome (PDB-5G06, EMDB-3366, reported resolution of 4.2 Å) [39]. In the first example of PDB-6M6Z, both Chain B and Chain D exhibited significant RSCC improvements compared to the deposited maps, increasing from 0.706 to 0.768 and from 0.739 to 0.777, respectively. In addition, 85.2% of residues in Chain B and 70.9% of residues in Chain D showed an increase in their RSCC scores. In the second example of PDB-5G06, the average RSCC scores increased from 0.601 to 0.635 for Chain B, with 82.9% of 240 residues showing improvement; and increased from 0.581 to 0.615 for Chain E, with 76.2% of 265 residues perform better.
Table A1: List of all EMDB/PDB examples in training sets.
Table A2: List of all EMDB/PDB examples in validation sets.
Table A3: List of all EMDB/PDB examples in test sets.
Table A4: List of all EMDB/PDB examples for protein structure modeling.

### v2

@@ 3.1 Dataset of protein structures and density maps
Our dataset was built with a set of cryo-EM density maps at resolutions from 4.0 Å to 7.9 Å from the EMDB databank [24] and their associated protein structures from the PDB databank [25]. To ensure that density maps are properly aligned with their corresponding PDB structures, we excluded maps and PDBs from the dataset if: (i) maps contain extensive regions without or misaligned corresponding PDB structures; (ii) maps contain other macromolecules except proteins; (iii) PDB structures contain backbone atoms only and/or unknown residues. Furthermore, to enhance training efficiency, we measured the correlation between map-PDB pairs using ChimeraX [26], and removed pairs with correlation score lower than 0.65. To avoid data redundancy, we measured the sequence identity between PDB structures, and retained only one if identity is greater than 30 %. As a result, a total of 384 pairs of cryo-EM maps and associated PDB structures remained. Among these data, 247 ($\sim$65 %), 62 ($\sim$15 %), and 75 ($\sim$20 %) map-PDB pairs were selected as training, validation, and test sets, respectively. Details are listed in Supplementary Tables S1 to S3.
@@ 4.1 Visualization and quantification of map enhancement
As displayed in Figure 2b, CryoSAMU significantly suppressed noise in the lip nanodisc regions (highlighted by dashed boxes in Figure 2a) of the deposited map for EMDB-24496 (PDB-7RKF). Moreover, the deposited map at a smaller volumes missed certain structural regions corresponding to the protein structures, as highlighted by black boxes in Figure 2a. In contrast, the CryoSAMU-enhanced maps exhibited better alignment with the corresponding protein structures, revealing more structural details, as demonstrated by black boxes in Figure 2b. Similar visual results were observed for another protein structure (see Supplementary Figure S1). Furthermore, residue-level real-space correlation coefficient (RSCC) measurements [35] in Figure 2c suggested significant improvements. Specifically, Chains A, B, and C exhibit RSCC increases compared to the deposited map, with correlations rising from 0.835 to 0.860, 0.848 to 0.865, and 0.801 to 0.836, respectively. In addition, 84.9%, 73.5%, and 90.6% of resiudes in Chains A, B, and C, respectively, showcased higher RSCC scores. Consistent RSCC improvements were also observed in other samples (see Supplementary Figure S2).
@@ 4.2 Benchmark I: improvement of real and Fourier space correlations
CC_box $\uparrow$ 0.731 0.679 0.618 0.862 0.834 0.751
CC_peaks $\uparrow$ 0.750 0.722 0.611 0.774 0.753 0.698
CC_volume $\uparrow$ 0.594 0.542 0.534 0.729 0.691 0.571
FSC05 $\downarrow$ 6.124 6.147 5.283 4.668 5.108 6.434
@@ 4.3 Benchmark II: improvement of protein structure modeling
As the goal of enhancing cryo-EM density maps is to improve the performance of protein structure modeling from density maps (i.e., map interpretability), we benchmarked protein structures constructed from CryoSAMU-enhanced maps against those processed by other methods. Specifically, we used a standard structure modeling tool, known as phenix.map_to_model [36], to construct protein structures from 20 maps enhanced by the different tested methods. These maps were randomly selected from the test dataset to ensure that they were not exposed during training, as listed in Supplementary Table S4. To evaluate these structures, we used phenix.chain_comparison [36] to compare the constructed structures against their corresponding ground-truth PDB protein structures. We reported two metrics: residue coverage and sequence match. The residue coverage indicates the fraction of residues in the query structure that match the corresponding residues in the target structure within 3.0 Å, regardless of residue type. The sequence match indicates the percentage of matched residues that share identical residue types between the query and target structures.
@@ 5 Conclusion and Discussion
In this work, we introduce CryoSAMU, the first structure-aware multimodal network for enhancing cryo-EM density maps at intermediate resolution of protein structures. Our approach combines 3D map features with corresponding structural features through cross-attention mechanisms. Notably, structural embeddings are used solely as auxiliary supervision during training to guide the model toward learning biologically-grounded and structure-aware representations from the density maps. This strategy aligns with the well-established principle of conditional training, where auxiliary modalities, despite not available at inference, are used during training to improve the quality and generalization of learned representations from the primary input [37]. In addition, we develop a self-attention weighting algorithm to produce fixed-size representations of structural embeddings derived from the pretrained ESM-IF1 model, preserving inter-chain and residue relationships while maintaining structural integrity. Our benchmark results demonstrate that CryoSAMU preforms competitively with existing cutting-edge methods, closely approaching the performance of EMReady, the current leading tool for cryo-EM density map enhancement. Notably, CryoSAMU achieves the fastest processing speed among all tested methods, positioning it as a promising solution for large-scale and practical applications. Furthermore, our ablation study reveals that incorporating an additional structural modality significantly boosts CryoSAMU’s performance across all evaluation metrics, suggesting a new avenue for future cryo-EM research to explore the effective integration of multimodal data during network training.
@@ Acknowledgments
Figure S1: Visualizations of deposited (blue) and CryoSAMU-enhanced (green) maps. The corresponding PDB structures (brown) are superimposed on the maps. a Human Dispatched-1 (PDB-6XE6, EMDB-22144, reported resolution of 4.53 Å) [38]. a-b: Maps displayed at the recommended contour level. c-d: Maps displayed at a higher contour level. Visualizations were produced by UCSF ChimeraX [26]. The protein structure modeling completeness and accuracy improved after CryoSAMU enhancement. For instance, residue coverage increased from 53.5% to 64.5%, as well as sequence match increased from 6.1% to 7.7%.
Figure S2: The real-space correlation coefficient (RSCC) comparison between deposited and CryoSAMU-enhanced maps. Top: a transmembrane nanopore TMH4C4 (PDB-6M6Z, EMDB-30126, reported resolution of 5.9 Å) [39]. Bottom: a yeast cytoplasmic exosome (PDB-5G06, EMDB-3366, reported resolution of 4.2 Å) [40]. In the first example of PDB-6M6Z, both Chain B and Chain D exhibited significant RSCC improvements compared to the deposited maps, increasing from 0.706 to 0.768 and from 0.739 to 0.777, respectively. In addition, 85.2% of residues in Chain B and 70.9% of residues in Chain D showed an increase in their RSCC scores. In the second example of PDB-5G06, the average RSCC scores increased from 0.601 to 0.635 for Chain B, with 82.9% of 240 residues showing improvement; and increased from 0.581 to 0.615 for Chain E, with 76.2% of 265 residues perform better.
Table S1: List of all EMDB/PDB examples in training sets.
Table S2: List of all EMDB/PDB examples in validation sets.
Table S3: List of all EMDB/PDB examples in test sets.
Table S4: List of all EMDB/PDB examples for protein structure modeling.